Hello, have a good day. I need quite some help here :(
The problem is of the "Concepts review problems" type. It challenges the student to find out tricks or shortcuts to work out problems without having specific numbers or values. I get always stuck here and I would like an explanation on how I should tackle the problem. How should I approach this type of questions:
"Let x and y be real numbers. If y > 1, compare: A. x+1/y+1 B. x/y +1"
Possible answers are: "A" if A is greater. "B" if B is greater. "C" if A is equal to B. "D" if we are not given enough information to work out an answer.
Thank you in advance.

Answer by Edwin McCravy(20081)

We are not given enough information. Only that y > 1. We are not given 
that x is not negative.  So here are three cases, one for A < B, one 
for A = B, and one for A > B.


 x   y         A=(x+1)/(y+1)                  B=(x/y)+1         A ? B
---------------------------------------------------------------------
 0   4    (0+1)/(4+1)=1/5=.2              0/4+1 = 0+1 = 1       A < B             
-4   2   (-4+1)/(2+1)=(-3)/3=-1      (-4)/(2)+1 = -2+1 =-1      A = B
-10  2   (-10+1)(2+1)=(-9)/3=-3       (-10)/2+1 = -4            A > B           

Incidentally if we had been given that x ≥ 0, then we could have 
concluded that A < B, but we can't since all we are given is that
x is a real number.

Edwin